a = nt + b / n (the right side is divided by n).


an = n^2 * tn + bn

Not sure. Non-homework.

Answer by rothauserc(4718)

(the right side is divided by n)
:
I assume you mean the right side of the =, then we have
:
a = (nt + b) / n
:
an = nt + b
:
assuming n is not = 0
:
If you meant the right side of the + sign, then
:
a = nt + (b/n)
:
an = (n^2)t + b
:
(n^2)t -an +b = 0
: